. Find the square of the following numbers.  iv. 93  
,jg,kymtdjrdtxhyrfjkhyktjuryhrdjyk,y

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


iv. (93)2
= (90+3 )2
= (90)2+ (3)2+ 2×90×3 =8100+9+540
= 8649



Thanks